OpenStudy (anonymous):

1. Whats the exact value of cot[cot^-1 sqrt3/3] ? ( is it infinity? ) 2. Whats the exact value of sin^-1(cos pi/2) ? ( is it 0? ) 3. Whats the exact value of tan(sin^-1 x/2) ?

OpenStudy (anonymous):

for the first one, at what angle is cotangent sqrt(3)/3

OpenStudy (anonymous):

do you not understand your own homework?

OpenStudy (anonymous):

at what angle is cotanget sqrt(3)/3?

OpenStudy (anonymous):

Nope, thats why I need help... I don't understand it 60 degrees?

OpenStudy (rogue):

1. The arccot of sqrt(3)/3 is 60 degrees, or pi/3. The cot of pi/3 is once again sqrt(3)/3. 2. Cosine of pi/2 is cosine of 90 degrees. The x-value at 90 degrees is 0, so cos pi/2 is 0. Arcsine of 0 is 0.

OpenStudy (rogue):

tan(sin^-1 x/2) Arcsine of x/2 will give you a certain angle, theta. We can draw a right triangle off this. Sine is opposite over hypotenuse.Tangent is opposite over adjacent. So tangent of theta is \[\tan \theta = \frac {x}{\sqrt {4-x^2}}\]